# Identifying Quantum Mechanical Statistics in Italian Corpora

### v1

@@ 1 Introduction
We have recently investigated the statistical behaviour of the words contained in medium- and large-size texts produced by English language, discovering that they obey Bose–Einstein, rather than the Maxwell–Boltzmann, statistics (Aerts & Beltran,, 2020; Beltran,, 2021; Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ; Aerts et al.,, 2023, 2024). We have also put forward the hypothesis that ‘meaning plays in the linguistic-conceptual domain a role that is very similar to the role played by quantum coherence in the physical domain (Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ; Aerts et al.,, 2024). This research fits a bold and novel research programme, called ‘quantum cognition’, which applies the mathematical formalism of quantum mechanics in Hilbert space to the modelling of high-level cognitive phenomena, e.g., categorization, language, perception, judgement, and decision-making (see, e.g., the volumes Van Rijsbergen, (2004); Khrennikov, (2010); Busemeyer & Bruza, (2012); Haven & Khrennikov, (2013); Melucci, (2015); Haven, Khrennikov & Robinson, (2017) and references therein). And, we have substantially contributed to the development of the quantum cognition programme (see, e.g., Aerts & Aerts, (2015); Aerts, (2009); Aerts et al., (2013); Aerts, Gabora & Sozzo, (2013); Aerts, Sassoli de Bianchi & Sozzo, (2016); Aerts & Sozzo, (2016); Aerts et al., (2018); Pisano & Sozzo, (2020); Aerts Arguëlles & Sozzo, (2020); Aerts et al., (2021)).
@@ 2 A theoretical framework for the statistical distribution of the words in a text
In the theoretical linguistic-conceptual framework we put forward, ‘words’ are names, or labels, for ‘concepts’. A concept is an abstract entity that we assume to be always in a defined ‘state’. The state of a concept captures the ‘meaning’ carried by the concept and can change under the influence of a ‘context’. Individual concepts can then be combined, or composed, to form more complex conceptual entities, as ‘conceptual combinations’ . Concepts and their combinations are examples of general ‘entities of meaning’. Next, a ‘text’ produced by human language, e.g., a story-telling text, is a complex entity of meaning made up of the concepts corresponding to all the words that appear in the text.
@@
We have recently studied a medium-size text, the story “In Which Piglet Meets a Heffalump” taken from Milne’s “Winnie the Pooh” (1926). The results are summarized in Tab. 1. We do not report complete details of the analysis, but limit ourselves to sketch the results here (the interested reader can make reference to Aerts & Beltran, (2020); Aerts & Beltran, 2022a ; Aerts & Beltran, 2022b ; Beltran, (2021); Aerts et al., (2023, 2024)). The Bose–Einstein distribution provides an almost perfect fit of the experimental numbers of appearance $N(E_{i})$, while the Maxwell–Boltzmann distribution does not provide a good fit at all. Hence, the “Winnie the Pooh” story statistically behaves as an ideal gas of bosons, which justifies the term “cogniton”, meant as the fundamental quantum of human language. We have also recently studied two large-size texts, namely, Well’s “The Magic Shop” (1903) and Swift’s “Gulliver’s Travels” (1726), obtaining exactly the same results, which are also reported in Tab. 1.
@@
The three aspects above, namely, presence of Bose–Einstein statistics, lack of statistical independence and incomplete indistiguishability, that we have identified in texts of English language, will be found back and widely extended when dealing with texts of Italian language, as we intend to prove in Sect. 3.666Though not the subject of the present article, we remind that the quantum theoretical framework illustrated in this section also provides a theoretical foundation of ‘Zipf’s law’ (Zipf,, 1935, 1949), which is systematically present in data across a range of cultural areas, even outside linguistic domains, and whose theoretical origin is currently unknown. Zipf’s law holds when Italian language is used too (Tuzzi, Popescu & Altman ,, 2009).
@@ 3 An empirical study through literary corpora of Italian language
Let us apply the quantum theoretical framework elaborated in Sect. 2 to the text “Cuore” (“Heart”). The most frequent word was Di, which was then attributed an energy level $E_{0}=0$, with a number of appearance $N(E_{0})=2,339$, followed by the word Che, which was attributed an energy level $E_{1}=1$, with a number of appearance $N(E_{1})=2,194$. Going down, the 20-th ranked word Father was attributed an energy level $E_{19}=19$, with a number of appearance $N(E_{19})=485$. Hence, the total energy of the word Father was $E_{19}N(E_{19})=19\times 485=9,215$, and so on.
@@
Firstly, we known from Sect. 1 that the appearance of a quantum statistical behaviour is attributed to the complete indistinguishability of identical entities in quantum mechanics, which is formally incorporated in the symmetry exchange postulate. However, some authors, including ourselves, maintain that this postulate is problematical (see also Sect. 6). On the other side, we have seen in Sect. 2 that identical words in a text are indistinguishable, while different words are not indistinguishable. We remind that identical words correspond to the same energy level, hence the same micro-state, of the cogniton, while different words generally correspond to different energy levels, hence different micro-states, of the cogniton. Let us take the example of the the word Romanzo (meaning Novel in English) in the text “Del Romanzo Storico” (“On the Historical Novel”), which has an energy level $E_{55}=55$ and a number of appearance $N(E_{55})=50$.
@@
It is evident that if we take two places where the word Romanzo appears, and exchange the two words, the overall meaning of the text does not change. This is because Romanzo, as it emerges as a concept in the meaning that the story holds, is an abstract entity within the story, i.e. a concept and not an object, and therefore identical and indistinguishable from any other word Romanzo that appears in the story. We believe that the appearance of quantum structures in large texts is partly due to the fact that a text that tells a story is primarily an entity of meaning, hence the words in it are conceptual and not objectual. Secondly, let us consider the word Storia (meaning History in English) which has an energy level $E_{18}=18$ and a number of appearance $N(E_{18})=137$ in the text “Del Romanzo Storico”, and consider again the word Romanzo. It is evident that the two words Romanzo and Storia, hence the corresponding micro-states, are not statistically independent, in the sense specified in Sects. 1 and 2. Indeed, suppose that some additional paragraphs are now added to the text. Since Storia occurs 137 times, while Romanzo occurs $50$ times, we expect that the probability that Storia occurs in the additional paragraphs is higher than the probability that Romanzo occurs in them. This is because the entire text carries an overall meaning, hence the words whose meaning is closer to that overall meaning have a higher chance to appear if the text is continued to be written. Using the terminology of physics, the more cognitons are in a given micro-state, the more a new cogniton is likely to be in that micro-state. Equivalently, the more a word is used in a text, the more likely it is that the same word will appear in a new piece of text based on the previous one. This expresses the typical tendency of bosons to clump, or cluster, together in the same micro-states. We believe that this lack of statistical independence in non-identical, hence not indistinguishable, words, is partly responsible for the appearance of Bose–Einstein statistics in the statistical behaviour of large texts.
@@ 4 The effect of word randomization
Let us consider again the five literary texts of Italian language analysed in Sect. 3 and investigate how the introduction of ‘randomness’ may affect the statistical distribution of the words appearing in the texts. The underlying idea is to partially remove the meaning of a text and check whether and how this removal will impact the statistics distribution of its words. Let us focus on the text “Senilità” (“As a Man Grows Older”), for the sake of brevity. We stress, however, that the results obtained on this text generally hold for all the others too. The text “Senilità” contains $n=9,278$ different words, hence $9,278$ energy levels, and a total number of words $N=63,082$ (see Tab. 2). Let us introduce the notion of ‘placeholder’, which will simplify our considerations. In the text “Senilità”, which we will call the ‘original text’ from now on, each word is in a given placeholder, thus we have $N=63,082$ placeholders overall. But, it is clear that a given word of the original text appears in various different placeholders, depending on its number of appearance in that text.
@@
We have used the tools offered by the website www.random.org to change the relative frequencies with which the different words of the “Senilità” text occupy the available 63,082 placeholders, determining new numbers of appearance in a genuinely random way. [^7] We note that the method of randomization developed in the website is linked to weather and atmospheric noise, thus it involves gases at room temperature. This is relevant to our scopes, because we will see in this and the following sections that word randomization can be considered as the linguistic analogue of a temperature increase. This operation of word randomization generates a new text with the same number $n=9,278$ of different words, hence again $9,278$ energy levels, and the same total number of words $N=63,082$, as the original text. We will call ‘randomized text’ from now on the new version of “Senilità” generated in this way. It should be noticed that this operation of word randomization is a subtle reshuffling of the relative frequencies of appearance of words, as it will become clear in the following.
@@
Let us preliminary state that this process of word randomization has a profound impact on the energy level carried by each word. Indeed, in the randomized text, the most frequent word has a number of appearance equal to 18, the second most frequent word has a number of appearance equal to 17, and so on. Let us now apply the quantum theoretical framework in Sect. 2 and assign again energy levels according to the decreasing number of appearance. In this way, we have generated a randomized text that has the same energy levels $E_{0}$, $E_{1}$, …, $E_{n-1}$ as the original text but different numbers of appearance $N(E_{0})$, $N(E_{1})$, …, $N(E_{n-1})$, i.e. different occupation numbers of these energy levels, which now correspond to a lower occupation of the lower energy levels. [^8] More specifically, regarding the text “Senilità”, we generated through the above mentioned website an amount $N=63,082$ of numbers taken out of the $n=9,278$ numbers. We made different trials and found that, while the most frequent number was different in each trial, the numbers of appearance were reasonably stable across trials. This operation was repeated on all texts discussed in Sect. 3. It is also important to notice that we have not entirely removed the meaning of the original text, because each word appearing in the randomized text has its own meaning and the randomized text still contains the same words and with the same total number as the original text. However, the numbers of appearance of each of the different words in the original text are now randomly changed, which entails that we have removed some of the meaning generated when the original text was written. We believe that the procedure described above has far reaching consequences on the development of a thermodynamic theory of language. Some of these consequences will be examined in detail in a forthcoming article (Aerts et al.,, 2025). But, we can already at this stage draw some interesting conclusions, as follows.
@@
In its original (non-randomized) version, “Senilità” showed the statistical behaviour illustrated in Figs. 1-3, hence the words in it obey Bose–Einstein statistics and show large deviations from Maxwell–Boltzmann statistics. We may pthen wonder whether the statistical distribution of the words contained in the randomized version can still be modelled by the Bose–Einstein distribution and, more important, whether the Maxwell–Boltzmann distribution still does not provide a good model at all. To this end, we apply the quantum theoretical framework presented in Sect. 2 to the randomized version of the text “Senilità” and determine the statistical distribution for the numbers of appearance $N(E_{i})$ that best fit the new empirical data, computing the pairs $(C,D)$ in Eq. (5) and $(A,B)$ of Eq. (5) subject to the conditions that the total number of words has to be $N=63,082$, as in the original text, while their total energy has now risen to $E=229,919,018$ (it was $E=62,516,178$ in the original text).
@@
To understand the result above, let us consider Tab.  3, where we report the total energy of words $E$, together with the fugacity $f=e^{\frac{\mu}{kT}}=\frac{1}{A}$ and the chemical potential $\mu=-B\ln A$ (see Sect. 2) of both the original and randomized versions of the text “Senilità”. If we look at the data, we realize at once the effect of word randomization. Indeed, we notice that, in the “Senilità” case, there has been an increase in the total energy $E$ of more than three times (from $62,516,178$ to $229,919,018$, see above). More important, the chemical potential $\mu$ has changed from $-5.0097$ to $-3,825.3644$. This is mainly due to the fact that the constant $B$ in Eq. (5) has drastically changed from $8,825.542$ to $54,902.4913$, i.e. more than six times. In physics, the constant $B$ expresses the temperature in energy units or, better, the heat energy. In other words, randomizing the words of the original text has dramatically increased the ‘heat’ of the text. These results suggest that, in the “Senilità” case, the effect produced by word randomization can be considered as a possible linguistic-conceptual counterpart of the effect produced by an increase of the temperature in a physical system.
@@
However, together with the notion of temperature, also the physical notion of ‘excitation’ should be introduced. Indeed, let us consider the most frequent word in the text “Senilità”, namely, Di. Its number of appearance is equal to $2,114$ in the original version of the text, while it drastically decreases to a value of $18$ in the randomized version. [^10] We are here assuming that the rank of the different words, in the transition between the original text and the randomized text, is preserved, similar to what would happen in a physical system, where the transfer of heat only leads to a change in the occupation numbers of the different energy levels, without changing the latter. This decrease can be interpreted as the linguistic-conceptual counterpart of a phenomenon of local excitation, i.e. transfer of kinetic energy provoked by heating, like the random bombardment provoked by a heath bath at room temperature (see also footnote 7). This local random effect entails that some of the words of the original text get excited, in the sense that they jump to a higher energy level, while others remain in the same energy level. In particular, out of $2,114$ placeholders of the word Di that were originally in the lowest energy level, only $18$ remain at this level, whereas the others get excited and jump to a higher energy level. And the same systematically happens with the other words of the text. In other words, we can describe the word randomization process also as a ‘word excitation’ process.

### v2

@@ 1 Introduction
We have recently investigated the statistical behaviour of the words contained in medium- and large-size texts produced by English language, discovering that they obey Bose–Einstein, rather than the Maxwell–Boltzmann, statistics (Aerts & Beltran,, 2020; Beltran,, 2021; Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ; Aerts et al., 2023a, ; Aerts et al.,, 2024). We have also put forward the hypothesis that ‘meaning plays in the linguistic-conceptual domain a role that is very similar to the role played by quantum coherence in the physical domain (Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ; Aerts et al.,, 2024). This research fits a bold and novel research programme, called ‘quantum cognition’, which applies the mathematical formalism of quantum mechanics in Hilbert space to the modelling of high-level cognitive phenomena, e.g., categorization, language, perception, judgement, and decision-making (see, e.g., the volumes Van Rijsbergen, (2004); Khrennikov, (2010); Busemeyer & Bruza, (2012); Haven & Khrennikov, (2013); Melucci, (2015); Haven, Khrennikov & Robinson, (2017) and references therein). And, we have substantially contributed to the development of the quantum cognition programme (see, e.g., Aerts & Aerts, (2015); Aerts, (2009); Aerts et al., (2013); Aerts, Gabora & Sozzo, (2013); Aerts, Sassoli de Bianchi & Sozzo, (2016); Aerts & Sozzo, (2016); Aerts et al., (2018); Pisano & Sozzo, (2020); Aerts Arguëlles & Sozzo, (2020); Aerts et al., (2021)).
@@ 2 A theoretical framework for the statistical distribution of the words in a text
In the theoretical linguistic-conceptual framework we put forward, ‘words’ are names, or labels, for ‘concepts’. A concept is an abstract entity that we assume to be always in a defined ‘state’. [^4] The introduction of concepts as abstract entities may suggest some analogies with other philosophical theories of concepts (see, e.g., Peacocke, (1992)). We stress, however, that we do not consider the terms ‘concept’ and ‘meaning’ as synonyms in our framework. We can best note that this is the case in human language by considering the combination of different concepts and noting how the meaning of such a combination usually comes about only when enough concepts of the combination have been reified and often even before they have all been reified. In this sense, ‘meaning’ is a notion that is carried in complex ways and potentially rather than already actually by individual concepts, and it is the overall mechanism that takes place during the combination of concepts, including the entanglement that occurs, that produces meaning. In our comparison of human language with quantum structures, we place meaning in terms of human language where quantum coherence is in terms of interacting quantum entities. Elaborating further on these aspects would distract us too far from the topic of this article and we refer the interested reader to our past publications (see, e.g., Aerts, Sassoli de Bianchi & Sozzo, (2016); Aerts et al., 2023b ; Aerts et al., 2024a ; Aerts et al., 2024b ). The state of a concept captures the ‘meaning’ carried by the concept and can change under the influence of a ‘context’. Individual concepts can then be combined, or composed, to form more complex conceptual entities, as ‘conceptual combinations’. Concepts and their combinations are examples of general ‘entities of meaning’. Next, a ‘text’ produced by human language, e.g., a story-telling text, is a complex entity of meaning made up of the concepts corresponding to all the words that appear in the text.
@@
We have recently studied a medium-size text, the story “In Which Piglet Meets a Heffalump” taken from Milne’s “Winnie the Pooh” (1926). The results are summarized in Tab. 1. We do not report complete details of the analysis, but limit ourselves to sketch the results here (the interested reader can make reference to Aerts & Beltran, (2020); Aerts & Beltran, 2022a ; Aerts & Beltran, 2022b ; Beltran, (2021); Aerts et al., 2023a ; Aerts et al., (2024)). The Bose–Einstein distribution provides an almost perfect fit of the experimental numbers of appearance $N(E_{i})$, while the Maxwell–Boltzmann distribution does not provide a good fit at all. Hence, the “Winnie the Pooh” story statistically behaves as an ideal gas of bosons, which justifies the term “cogniton”, meant as the fundamental quantum of human language. We have also recently studied two large-size texts, namely, Well’s “The Magic Shop” (1903) and Swift’s “Gulliver’s Travels” (1726), obtaining exactly the same results, which are also reported in Tab. 1.
@@
The three aspects above, namely, presence of Bose–Einstein statistics, lack of statistical independence and incomplete indistiguishability, that we have identified in texts of English language, will be found back and widely extended when dealing with texts of Italian language, as we intend to prove in Sect. 3.777Though not the subject of the present article, we remind that the quantum theoretical framework illustrated in this section also provides a theoretical foundation of ‘Zipf’s law’ (Zipf,, 1935, 1949), which is systematically present in data across a range of cultural areas, even outside linguistic domains, and whose theoretical origin is currently unknown. Zipf’s law holds when Italian language is used too (Tuzzi, Popescu & Altman ,, 2009).
@@ 3 An empirical study through literary corpora of Italian language
Let us apply the quantum theoretical framework elaborated in Sect. 2 to the text “Cuore” (“Heart”). The most frequent word was Di, which was then attributed an energy level $E_{0}=0$, with a number of appearance $N(E_{0})=2,180$, followed by the word Che, which was attributed an energy level $E_{1}=1$, with a number of appearance $N(E_{1})=2,071$. Going down, the 22nd ranked word Father was attributed an energy level $E_{21}=21$, with a number of appearance $N(E_{21})=426$. Hence, the total energy of the word Father was $E_{21}N(E_{21})=21\times 426=8,946$, and so on.
@@
Firstly, we known from Sect. 1 that the appearance of a quantum statistical behaviour is attributed to the complete indistinguishability of identical entities in quantum mechanics, which is formally incorporated in the symmetry exchange postulate. However, some authors, including ourselves, maintain that this postulate is problematical (see also Sect. 6). On the other side, we have seen in Sect. 2 that identical words in a text are indistinguishable, while different words are not indistinguishable. We remind that identical words correspond to the same energy level, hence the same micro-state, of the cogniton, while different words generally correspond to different energy levels, hence different micro-states, of the cogniton. Let us take the example of the the word Romanzo (meaning Novel in English) in the text “Del Romanzo Storico” (“On the Historical Novel”), which has an energy level $E_{51}=51$ and a number of appearance $N(E_{51})=52$.
@@
It is evident that if we take two places where the word Romanzo appears, and exchange the two words, the overall meaning of the text does not change. This is because Romanzo, as it emerges as a concept in the meaning that the story holds, is an abstract entity within the story, i.e. a concept and not an object, and therefore identical and indistinguishable from any other word Romanzo that appears in the story. We believe that the appearance of quantum structures in large texts is partly due to the fact that a text that tells a story is primarily an entity of meaning, hence the words in it are conceptual and not objectual. Secondly, let us consider the word Storia (meaning History in English) which has an energy level $E_{18}=18$ and a number of appearance $N(E_{18})=139$ in the text “Del Romanzo Storico”, and consider again the word Romanzo. It is evident that the two words Romanzo and Storia, hence the corresponding micro-states, are not statistically independent, in the sense specified in Sects. 1 and 2. Indeed, suppose that some additional paragraphs are now added to the text. Since Storia occurs $139$ times, while Romanzo occurs $52$ times, we expect that the probability that Storia occurs in the additional paragraphs is higher than the probability that Romanzo occurs in them. This is because the entire text carries an overall meaning, hence the words whose meaning is closer to that overall meaning have a higher chance to appear if the text is continued to be written. Using the terminology of physics, the more cognitons are in a given micro-state, the more a new cogniton is likely to be in that micro-state. Equivalently, the more a word is used in a text, the more likely it is that the same word will appear in a new piece of text based on the previous one. This expresses the typical tendency of bosons to clump, or cluster, together in the same micro-states. We believe that this lack of statistical independence in non-identical, hence not indistinguishable, words, is partly responsible for the appearance of Bose–Einstein statistics in the statistical behaviour of large texts.
@@ 4 The effect of word randomization
Let us consider again the five literary texts of Italian language analysed in Sect. 3 and investigate how the introduction of ‘randomness’ may affect the statistical distribution of the words appearing in the texts. The underlying idea is to partially remove the meaning of a text and check whether and how this removal will impact the statistics distribution of its words. Let us focus on the text “Senilità” (“As a Man Grows Older”), for the sake of brevity. We stress, however, that the results obtained on this text generally hold for all the others too. The text “Senilità” contains $n=9,024$ different words, hence $9,024$ energy levels, and a total number of words $N=63,339$ (see Tab. 2). Let us introduce the notion of ‘placeholder’, which will simplify our considerations. In the text “Senilità”, which we will call the ‘original text’ from now on, each word is in a given placeholder, thus we have $N=63,082$ placeholders overall. But, it is clear that a given word of the original text appears in various different placeholders, depending on its number of appearance in that text.
@@
We have used the tools offered by the website www.random.org to change the relative frequencies with which the different words of the “Senilità” text occupy the available $63,339$ placeholders, determining new numbers of appearance in a genuinely random way. [^8] We note that the method of randomization developed in the website is linked to weather and atmospheric noise, thus it involves gases at room temperature. This is relevant to our scopes, because we will see in this and the following sections that word randomization can be considered as the linguistic analogue of a temperature increase. This operation of word randomization generates a new text with the same number $n=9,024$ of different words, hence again $9,024$ energy levels, and the same total number of words $N=63,339$, as the original text. We will call ‘randomized text’ from now on the new version of “Senilità” generated in this way. It should be noticed that this operation of word randomization is a subtle reshuffling of the relative frequencies of appearance of words, as it will become clear in the following.
@@
Let us preliminary state that this process of word randomization has a profound impact on the energy level carried by each word. Indeed, in the randomized text, the most frequent word has a number of appearance equal to $21$, the second most frequent word has a number of appearance equal to $19$, and so on. Let us now apply the quantum theoretical framework in Sect. 2 and assign again energy levels according to the decreasing number of appearance. In this way, we have generated a randomized text that has the same energy levels $E_{0}$, $E_{1}$, …, $E_{n-1}$ as the original text but different numbers of appearance $N(E_{0})$, $N(E_{1})$, …, $N(E_{n-1})$, i.e. different occupation numbers of these energy levels, which now correspond to a lower occupation of the lower energy levels. [^9] More specifically, regarding the text “Senilità”, we generated through the above mentioned website an amount $N=63,339$ of numbers taken out of the $n=9,024$ numbers. We made different trials and found that, while the most frequent number was different in each trial, the numbers of appearance were reasonably stable across trials. This operation was repeated on all texts discussed in Sect. 3. It is also important to notice that we have not entirely removed the meaning of the original text, because each word appearing in the randomized text has its own meaning and the randomized text still contains the same words and with the same total number as the original text. However, the numbers of appearance of each of the different words in the original text are now randomly changed, which entails that we have removed some of the meaning generated when the original text was written. We believe that the procedure described above has far reaching consequences on the development of a thermodynamic theory of language. Some of these consequences will be examined in detail in a forthcoming article (Aerts et al.,, 2025). But, we can already at this stage draw some interesting conclusions, as follows.
@@
In its original (non-randomized) version, “Senilità” showed the statistical behaviour illustrated in Figs. 1-3, hence the words in it obey Bose–Einstein statistics and show large deviations from Maxwell–Boltzmann statistics. We may pthen wonder whether the statistical distribution of the words contained in the randomized version can still be modelled by the Bose–Einstein distribution and, more important, whether the Maxwell–Boltzmann distribution still does not provide a good model at all. To this end, we apply the quantum theoretical framework presented in Sect. 2 to the randomized version of the text “Senilità” and determine the statistical distribution for the numbers of appearance $N(E_{i})$ that best fit the new empirical data, computing the pairs $(C,D)$ in Eq. (5) and $(A,B)$ of Eq. (5) subject to the conditions that the total number of words has to be $N=63,339$, as in the original text, while their total energy has now risen to $E=225,613,236$ (it was $E=60,296,678$ in the original text).
@@
To understand the result above, let us consider Tab.  3, where we report the total energy of words $E$, together with the fugacity $f=e^{\frac{\mu}{kT}}=\frac{1}{A}$ and the chemical potential $\mu=-B\ln A$ (see Sect. 2) of both the original and randomized versions of the text “Senilità”. If we look at the data, we realize at once the effect of word randomization. Indeed, we notice that, in the “Senilità” case, there has been an increase in the total energy $E$ of more than three times (from $60,296,678$ to $225,613,236$, see above). More important, the chemical potential $\mu$ has changed from $-4.4073$ to $-3,829.26$. This is mainly due to the fact that the constant $B$ in Eq. (5) has drastically changed from $8,713.51$ to $55,935.53$, i.e. more than six times. In physics, the constant $B$ expresses the temperature in energy units or, better, the heat energy. In other words, randomizing the words of the original text has dramatically increased the ‘heat’ of the text. These results suggest that, in the “Senilità” case, the effect produced by word randomization can be considered as a possible linguistic-conceptual counterpart of the effect produced by an increase of the temperature in a physical system.
@@
However, together with the notion of temperature, also the physical notion of ‘excitation’ should be introduced. Indeed, let us consider the most frequent word in the text “Senilità”, namely, Di. Its number of appearance is equal to $2,120$ in the original version of the text, while it drastically decreases to a value of $21$ in the randomized version. [^11] We are here assuming that the rank of the different words, in the transition between the original text and the randomized text, is preserved, similar to what would happen in a physical system, where the transfer of heat only leads to a change in the occupation numbers of the different energy levels, without changing the latter. This decrease can be interpreted as the linguistic-conceptual counterpart of a phenomenon of local excitation, i.e. transfer of kinetic energy provoked by heating, like the random bombardment provoked by a heath bath at room temperature (see also footnote 8). This local random effect entails that some of the words of the original text get excited, in the sense that they jump to a higher energy level, while others remain in the same energy level. In particular, out of $2,120$ placeholders of the word Di that were originally in the lowest energy level, only $21$ remain at this level, whereas the others get excited and jump to a higher energy level. And the same systematically happens with the other words of the text. In other words, we can describe the word randomization process also as a ‘word excitation’ process.